OpenStudy (anonymous):

Given D (6, 6) and M (2,4) find E.

OpenStudy (anonymous):

E=42... IS THERE A DIAGRAM?

OpenStudy (anonymous):

\[\frac{x+6}{2}=2\] \[\frac{y+6}{2}=4\]

OpenStudy (anonymous):

no diagram

OpenStudy (anonymous):

solve these for x and y \[x+6=4\] \[x=-2\] \[y+6=8\] \[y=2\]

OpenStudy (anonymous):

answer is (-2,2)

OpenStudy (mathteacher1729):

Again, drawing should help make sense of the formulas... (see attached pic)

OpenStudy (mathteacher1729):

*edit: I wrote "D" instead of "B". The method is still the same though. :)